I can't seem to find the value of x in a ratio problem.
4/x = 7/10

Found 2 solutions by saifoo.khan, fcabanski:
Answer by saifoo.khan(162)

4/x = 7/10
Cross multiply,
7x = 4*10
7x = 40
x = 40/7

Answer. =)
If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($15/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($1.5-$3.0 per problem). Contact me at einstein.saifoo@gmail.com

Answer by fcabanski(1391)

4/x = 7/10


Cross multiply.
7x=40
x=40/7

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com